How many circles of radius 9 cm can be cut from a rectangular sheet 3.7 m by 1.05m? Thank you.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Along the long axis, 41 circles *9 cm=369 cm.
Along the short axis 11 circles can be cut from 105 cm
451 circles.

Answer by ikleyn(52776)

.
How many circles of radius 9 cm can be cut from a rectangular sheet 3.7 m by 1.05m? Thank you.
------------------------------------------------------------------------

The diameter of these circles is equal to 2*9 cm = 18 cm.

Therefore,
along the long axis, 20 circles * 18 cm = 360 cm.

Along the short axis 5 circles can be cut from 105 cm (5*18 = 90 cm)

In total, 20*5 = 100 circles.